# MARM: Unlocking the Future of Recommendation Systems through Memory Augmentation and Scalable Complexity

### v1

@@ Abstract.
Scaling-law has guided the language model designing for past years, to estimate expected model performance with respect to the size of learnable parameters and training samples data-scales, e.g. GPTs. It is worth noting that the scaling laws of NLP cannot be directly applied to recommendation systems due to the following reasons: (1) The amount of training samples and model parameters is typically not the bottleneck for the model. Our recommendation system can generate over 50 billion user samples daily, and such a massive amount of training data can easily allow our model parameters to exceed 200 billion, surpassing many LLMs (about 100B). (2) To ensure the stability and robustness of the recommendation system, it is essential to control computational complexity FLOPs carefully. In training, we need to process a vast number of recommendation samples every day. During online inference, we usually need to respond within milliseconds (LLMs are usually a few seconds). Considering the above differences with LLM, we can draw a conclusion that: for a RecSys model, compared to model parameters, the computational complexity FLOPs is the more expensive factor that requires careful control.
In this paper, we propose our milestone work, MARM (Memory Augmented Recommendation Model), which explores a new cache scaling-laws successfully. By caching part of complex module calculation results, Our MARM extends the single-layer attention-based sequences interests modeling module to a multiple-layer setting with minor inference complexity FLOPs cost (i.e, module time complexity $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$). Equipped with the cache idea, our MARM solution significantly overcomes computational bottlenecks, and could seamlessly empower all user sequences interest extracting modules, and even other models. To support our MARM, we construct a 60TB cache storage center for offline training and online serving. Comprehensive experiment results show that our MARM brings offline 0.43% GAUC improvements and online 2.079% play-time per user gains. Our MARM has been deployed on Kwai short-videos platform,, serving tens of millions of users daily.
@@ 1. Introduction
  • We give detailed reasons why the scaling-laws of NLP do not match the ranking scenario, and we are the first work to explore new scaling-laws under the wide-used ranking model architecture from a fresh cache-performance perspective.
@@ 2.3. MARM to Support Other Models
  • For Cascading model, we implement a same multi-layer target-attention process as our Ranking model in Eq.(6). It is important to note that in the cascading model, we do not perform re-accumulation of the MARM cache; instead, we directly use the already accumulated MARM cache. This approach saves resources while also enhancing the consistency between cascading and ranking models. However, due to limited computational power, we also only used latest short-term sequences of users, rather than GSU searched cache results for each item candidate.
@@ 3.2. Discussion of Cache Scaling Laws
Before going on, we first present an important concept, the cache size $\mathcal{C}=L*n*d$, where L is the attention layer depth of MARM, n is the length of the item sequence, and d is the representation dimension. It is worth noting that the computational complexity and storage size of the cache size are both linearly proportional, which ensures that the increase in FLOPs is offset by the efficiencies gained from the cache techniques.
@@ 3.2.3. Scaling up the cache size $\mathcal{C}$
Table 1. Offline comparison with SOTAs. Each module is added individually to the baseline model. The best and runner-up results are highlighted in bold/underlined.
AUC ($mean\pm std$ ) GAUC ($mean\pm std$)
Baseline $0.8233\pm 0.0001$ $0.7093\pm 0.00005$
DIN $0.8235\pm 0.00019$ $0.7010\pm 0.00012$
SIM Soft $0.8241\pm 0.00004$ $0.7109\pm 0.00007$
TWIN $0.8249\pm 0.00005$ $0.7142\pm 0.00009$
TWIN V2 $0.8251\pm 0.00023$ $0.7147\pm 0.00007$
MARM(L=1) $0.8256\pm 0.00008$ $0.7157\pm 0.00006$
MARM(L=2) $0.8262\pm 0.0004$ $0.7169\pm 0.00012$
MARM(L=4) $\mathbf{0.8270\pm 0.0003}$ $\mathbf{0.7183\pm 0.00008}$
Improvement $0.19\%$ $0.36\%$
@@ 3.3. Real-World Experiments
AUC ($mean\pm std$ ) GAUC ($mean\pm std$)
@@
DIN $0.8235\pm 0.00019$ $0.7010\pm 0.00012$
   +SIM Soft $0.8251\pm 0.00004$ $0.7111\pm 0.00008$
      +TWIN $0.8260\pm 0.00005$ $0.7156\pm 0.00006$
         +TWIN V2 $0.8268\pm 0.00023$ $0.7167\pm 0.00011$
            +HSTU* $0.8268\pm 0.00017$ $0.7168\pm 0.00021$
            +MARM(L=1) $0.8270\pm 0.00023$ $0.7180\pm 0.00020$
            +MARM(L=2) $0.8272\pm 0.00019$ $0.7191\pm 0.00021$
            +MARM(L=4) $\mathbf{0.8286\pm 0.00013}$ $\mathbf{0.7210\pm 0.00011}$
Improvement $0.22\%$ $0.43\%$
Figure 6. GSU Top-k overlapped rate between different layers.
Table 3. Online A/B testing results of Short-Video services at Kwai.
Stages Watch Time Metrics Interaction Metrics
Average Watch Time Watch Time Video View Like Comment Forward Follow
Retrieval +0.489% +0.456% +0.463% +0.167% +0.675% -0.550% +0.288%
Cascading +0.276% +0.260% +0.086% +$0$017% +0.495% +0.870%
Ranking +1.314% +1.370% +0.103% +0.605% -0.669% -0.114% -0.227%
@@ 3.3.1. Compared Approaches Details
  • HSTU*(Zhai et al., 2024): This method employs an HSTU-style multi-layer self-masked attention mechanism to model user history (time complexity $\mathcal{O}(n^{2}*d)$). Due to computational limitations, each historical item only uses a few features: item ID, author ID, tag, and user feedback, with a history length of 2000 and an attention depth of 4.
@@ 3.3.2. Individual Experiment Performance Comparisons (RQ1)
In this section, we introduce different sequence modeling modules into the baseline model, and the experimental results are shown in Table 1. MARM achieves the best performance compared to all the baselines. It is worth noting that although MARM does not use a user history length in the range of 100,000 like TWINV2, but only 6000, its performance still significantly outperforms TWINV2. Thanks to MARM’s ability to scale linearly in depth, we can expand the attention depth, not just the sequence length of the users. Furthermore, we can observe that with the increase in attention depth, MARM still shows significant improvement.
@@ 3.3.5. MARM Cost Discussion (RQ4)
Improving the attention depth in the user sequence modeling module of recommendation models is a very challenging task. HSTU follows a first sample-aggregation then learning paradigm to alleviate calculation pressure, its overall computational complexity for a user with a lifelong historical behavior is about $O(L*n^{2}*d*(\frac{N}{r}))$, where N denote the total interaction amount of a user, and $\textit{r}<<n$ is the aggregation level of user samples, meaning that every r user samples are aggregated together for training. In practice, a larger aggregation r typically reduces the modeling complexity but can lead to delays in user feedback entering the model training, and easily results in a decline in online metrics, creating a trade-off between cost and effectiveness.
@@ 3.3.6. Online Results (RQ5)
Notably, during the fine ranking stage, MARM achieved a significant increase in average user watch time of +1.314% and a watch time gain of +1.370%, highlighting its crucial role in enhancing user watch time. In the Retrieval stage, MARM also demonstrated remarkable gains, with average user watch time and total watch time increasing by +0.489% and +0.456%, respectively. While there were some negative impacts on Comment/Forward/Follow, significant gains were achieved in user interactions through Like, thereby the online results still within a reasonable range of substitution. Furthermore, in the Cascading ranking stage, MARM resulted in a notable increase in average user watch time of +0.276%, with overall interactions remaining largely positive. Overall, our online results demonstrate that our MARM not only could extend Ranking stage, but also enable both the Cascading and Retrieval stages to effectively leverage its computational capabilities.
@@ 4. Conclusion
We propose the Memory Augmented Recommendation Model (MARM), a multi-layer recommendation model that utilizes Memory Cache to accelerate inference. In recommendation scenarios, computational complexity is a significant performance constraint. We use caching to store partial computation results of complex models, reducing the complexity of single modeling from $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$. Based on the MARM framework, we can extend sequence modeling from single-layer target attention to multi-layer with linear resource consumption, significantly breaking through the computational bottleneck of recommendation models and enabling modeling of users’ life-long histories. We explored the scaling law within the MARM framework, confirming a proportional relationship between cache size and recommendation performance. Furthermore, our MARM method can seamlessly integrate with existing recommendation models, including fine ranking, coarse ranking, and recall. Through the MARM series of methods, we achieved 2.079% increase in average app viewing time per user in Kuaishou’s overseas operations, demonstrating the effectiveness and practicality of the MARM model. To the best of our knowledge, MARM is the first recommendation model to achieve multi-dimensional scaling in the field of life-long sequence modeling.

### v2

@@ Abstract.
Scaling-law has guided the language model design for past years, to estimate expected model performance with respect to the size of learnable parameters and training samples data-scales, e.g. GPTs. It is worth noting that the scaling laws of NLP cannot be directly applied to recommendation systems due to the following reasons: (1) The amount of training samples and model parameters is typically not the bottleneck for the model. Our recommendation system can generate over 50 billion user samples daily, and such a massive amount of training data can easily allow our model parameters to exceed 200 billion, surpassing many LLMs (about 100B). (2) It is essential to control computational complexity FLOPs carefully in recommendation system. In training, we need to process a vast number of recommendation samples every day. During online inference, we must respond within milliseconds (LLMs take usually a few seconds). Considering the above differences with LLM, we can conclude that: for a RecSys model, compared to model parameters, the computational complexity FLOPs is a more expensive factor that requires careful control.
In this paper, we propose our milestone work, MARM (Memory Augmented Recommendation Model), which explores a new cache scaling-laws successfully. By caching part of complex module calculation results, Our MARM extends the single-layer attention-based sequences interests modeling module to a multiple-layer setting with minor inference complexity FLOPs cost (i.e, module time complexity $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$). Equipped with the cache idea, our MARM solution significantly overcomes computational bottlenecks, and could seamlessly empower all user sequences interest extracting modules, and even other models. To support our MARM, we construct a 60TB cache storage center for offline training and online serving. Comprehensive experiment results show that our MARM brings offline 0.43% GAUC improvements and online 2.079% play-time per user gains. Our MARM has been deployed on Kwai short-videos platform, serving tens of millions of users daily.
@@ 1. Introduction
  • We are the first work to explore new scaling-laws under the wide-used ranking model architecture from a fresh cache-performance perspective.
@@ 2.1.2. Cache Memory Storage Look-Up
Figure 3. MARM realtime cache update&use workflow;The blue box illustrates the process of the first displayed item being updated into the cache, followed by the subsequent retrieval and usage from the cache.
@@ 2.3. MARM to Support Other Models
  • For Cascading model, we implement a same multi-layer target-attention process as our Ranking model in Eq.(6). It is important to note that in the cascading model, we do not perform re-accumulation of the MARM cache; instead, we directly use the already accumulated MARM cache and only used latest short-term sequences of users. This approach saves resources while also enhancing the consistency between cascading and ranking models.
@@ 3.2. Discussion of Cache Scaling Laws
Before going on, we first present an important concept, the cache size $\mathcal{C}=L*n*d$, where L is the attention layer depth of MARM, n is the length of the item sequence, and d is the representation dimension. It is worth noting that both the computational complexity of the cache size and the amount of storage resources are linearly proportional to $\mathcal{C}$. This means that the size of $\mathcal{C}$ is linearly positively correlated with the online latency of the recommendation system’s MARM module, as well as with all resources involved in training, inference, and storage, ensuring that the increase in FLOPs is offset by the efficiencies gained from the cache techniques.
@@ 3.3. Real-World Experiments
Table 1. Offline comparison with SOTAs. Each module is added individually to the baseline model. The arrows in the data on MARM represent a comparison of the FLOPS of the uncache version of MARM. The best and runner-up results are highlighted in bold/underlined.
AUC ($mean\pm std$ ) GAUC ($mean\pm std$) FLOPS
@@
DIN $0.8235\pm 0.00019$ $0.7010\pm 0.00012$ 3.3M
SIM Soft $0.8241\pm 0.00004$ $0.7109\pm 0.00007$ 4.6M
TWIN $0.8249\pm 0.00005$ $0.7142\pm 0.00009$ 500M
TWIN V2 $0.8251\pm 0.00023$ $0.7147\pm 0.00007$ 830M
MARM(L=1) $0.8256\pm 0.00008$ $0.7157\pm 0.00006$ 700M $\leftarrow$ 10.4B*
MARM(L=2) $0.8262\pm 0.0004$ $0.7169\pm 0.00012$ 900M $\leftarrow$ 20.6B*
MARM(L=4) $\mathbf{0.8270\pm 0.0003}$ $\mathbf{0.7183\pm 0.00008}$ 1.3B $\leftarrow$ 40.8B*
Improvement $0.19\%$ $0.36\%$
The FLOPS marked with * indicates the un-cache version MARM needed.
@@ 3.3.1. Compared Approaches Details
  • HSTU*(Zhai et al., 2024): This method employs an HSTU-style multi-layer self-masked attention mechanism to model user history (time complexity $\mathcal{O}(n^{2}*d)$). Due to computational limitations, each historical item only uses a few features: item ID, author ID, tag, and user feedback, with 2000 history length and 4 depth. Its FLOPs can be seen as the un-cache version MARM.
@@ 3.3.2. Individual Experiment Performance Comparisons (RQ1)
In this section, we introduce different sequence modeling modules into the baseline model, and the experimental results are shown in Table 1. MARM achieves the best performance compared to all the baselines. It is worth noting that although MARM does not use a user history length in the range of 100,000 like TWINV2, but only 6000, its performance still significantly outperforms TWINV2. We also list the FLOPS metrics for each module. It should be noted that in the two-stage modeling architecture of SIM, which consists of GSU and ESU, although the sequence length of the ESU during training is only a few hundred, for fairness, we include the computational cost of the previously calculated GSU part when calculating FLOPS. Thanks to MARM’s ability to scale linearly in depth, we can expand the attention depth, not just the sequence length of the users. Furthermore, we can observe that with the increase in attention depth, MARM still shows significant improvement.
AUC ($mean\pm std$ ) GAUC ($mean\pm std$) FLOPS
Baseline $0.8233\pm 0.0001$ $0.7093\pm 0.00005$
DIN $0.8235\pm 0.00019$ $0.7010\pm 0.00012$ 3.3M
   +SIM Soft $0.8251\pm 0.00004$ $0.7111\pm 0.00008$ 7.9M
      +TWIN $0.8260\pm 0.00005$ $0.7156\pm 0.00006$ 508M
         +TWIN V2 $0.8268\pm 0.00023$ $0.7167\pm 0.00011$ 838M
            +HSTU* $0.8268\pm 0.00017$ $0.7168\pm 0.00021$ 5.5B
[…]
Improvement $0.22\%$ $0.43\%$
Figure 6. GSU Top-k overlapped rate between different layers.
Table 3. Online A/B testing results of Short-Video services at Kwai.
Stages Retention&Watch Time Metrics Interaction Metrics
1D Retention 3D Retention Average Watch Time Watch Time Video View Like Comment Forward Follow
Retrieval +0.02% +0.009% +0.489% +0.456% +0.463% +0.167% +0.675% -0.550% +0.288%
Cascading +0.01% +0.01% +0.276% +0.260% +0.086% +$0$017% +0.495% +0.870%
Ranking +0.07% +0.13% +1.314% +1.370% +0.103% +0.605% -0.669% -0.114% -0.227%
@@ 3.3.5. MARM Cost Discussion (RQ4)
As an equivalent form of an un-cached MARM, HSTU follows a first sample-aggregation then learning paradigm to alleviate calculation pressure, its overall computational complexity for a user with a lifelong historical behavior is about $O(L*n^{2}*d*(\frac{N}{r}))$, where N denote the total interaction amount of a user, and $\textit{r}<<n$ is the aggregation level of user samples, meaning that every r user samples are aggregated together for training. In practice, a larger aggregation r typically reduces the modeling complexity but can lead to delays in user feedback entering the model training, and easily results in a decline in online metrics, creating a trade-off between cost and effectiveness.
@@ 3.3.6. Online Results (RQ5)
Our app’s retention metrics are positive, indicating that the launch of MARM has resulted in a better user experience.Notably, during multiple stage, MARM achieved a significant increase in average user watch time of +2.079%, highlighting its crucial role in enhancing user watch time. While there were slight negative impacts on Comment/Forward/Follow, significant gains were achieved in user interactions through Like. Since the primary goal of the recommendation system is to maximize average watch time, the online results still within a reasonable range of substitution.
@@ 4. Conclusion
We propose the Memory Augmented Recommendation Model (MARM), a multi-layer recommendation model that utilizes Memory Cache to accelerate inference. In recommendation scenarios, computational complexity is a significant performance constraint. We use caching to store partial computation results of complex models, reducing the complexity of single modeling from $\mathcal{O}(n^{2}*d)\rightarrow\mathcal{O}(n*d)$. Based on the MARM framework, we can extend sequence modeling from single-layer target attention to multi-layer with linear resource consumption, significantly breaking through the computational bottleneck of recommendation models and enabling modeling of users’ life-long histories. We explored the scaling law within the MARM framework, confirming a proportional relationship between cache size and recommendation performance. Furthermore, our MARM method can seamlessly integrate with existing recommendation models, including fine ranking, coarse ranking, and retrieval. Through the MARM series of methods, we achieved 2.079% increase in average app viewing time per user in Kuaishou’s overseas operations, demonstrating the effectiveness and practicality of the MARM model. To the best of our knowledge, MARM is the first recommendation model to achieve multi-dimensional scaling in the field of life-long sequence modeling.